if a system has an infinite number of solutions, use set-builder notation to write the solution set. If a system has no solution, state this.
9x − 2y = 3,
3x − 6 = y

Answer by fcabanski(1391)

Solve this with substitution. The second equation indicates that y=3x-6. Substitute that for y in the first equation.


9x-2(3x-6)=3 | Distribute the 2
9x-6x+12=3 | combine like terms and subtract 12 from both sides
3x=-9 | divide both sides by 3
x=-3


Substitute x=-3 into the second equation, to solve for y.
3(-3)-6=y
-9-6=y
-15=y


If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com